### v5

@@ 6.1. Good-for-MDP automata
We first define MDPs and good-for-MDP automata. A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$. Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions. Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $a\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $a$ is taken. Accordingly, for every $s\in S$ and $a\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})=1$. Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
@@
Consider a nondeterministic automaton $\mathcal{N}$} with alphabet $\Sigma$, and a $\Sigma$-labeled MDP $\mathcal{M}$}, and consider the following game between a player and nature. The positions of the game are pairs $\left\langle s,q\right\rangle$, where $s\in S$ is a state of $\mathcal{M}$}, and $q\in Q$ is a state of $\mathcal{N}$}. The game starts by the player picking an action $a\in A_{s_{0}}$ available from the initial state of $\mathcal{M}$}, and the game moves to $\left\langle s,q_{0}\right\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},a,s)$. Then, at each round from position $\left\langle s,q\right\rangle$, the player chooses an action $a\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$, and the game moves to $\left\langle s^{\prime},q^{\prime}\right\rangle$. The objective of the player is to on the one hand maximize the probability of generating a word in $L({\mathcal{N}})$ and on the other hand, to generate a rejecting run on words in $L({\mathcal{N}})$ with zero probability. When the automaton $\mathcal{N}$} is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique. When the automaton is nondeterministic, the player may fail to generate an accepting run of $\mathcal{N}$} even when the actions he provides induce a word in $L({\mathcal{N}})$. Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{N}})$. Intuitively, $\mathcal{N}$} is good-for-MDP, if for every MDP $\mathcal{M}$}, the nondeterminism in $\mathcal{N}$} does not prevent the player from maximizing the probability of generating a word in $L({\mathcal{N}})$ while generating a rejecting run on words in $L({\mathcal{N}})$ with zero probability. We now formalize this intuition.
@@
A run of $\mathcal{M}$} is an infinite sequence of states $r=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $a\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},a,s_{i+1})>0$. A finite run is a prefix of such a sequence. Let $\Omega({\mathcal{M}})\subseteq S^{\omega}$ denote the set of runs of $\mathcal{M}$} from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})\subseteq S^{*}$ denote the set of finite such runs. For a run $r\in\Omega({\mathcal{M}})$, we define the corresponding labeled run as $\tau(r)=\tau(s_{1}),\tau(s_{2}),\tau(s_{3}),\ldots\in\Sigma^{\omega}$. Note that the label of the first state $s_{0}$ of a run is not part of $\tau(r)$.
@@
Consider the set $A$ of actions. Let $\mathrm{dist}(A)$ be the set of distributions over $A$. A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $a\in A$ such that $d(a)=1$. A strategy for an MDP $\mathcal{M}$} is a function $\mu:\Pi({\mathcal{M}})\rightarrow\mathrm{dist}(A)$, which suggests to the player a distribution on the available actions given the history of the game so far. The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(a)>0$ implies that $a\in A_{s_{k}}$. Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$. Formally, we have that $\left\langle s_{0}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\left\langle s_{0},s_{1},\ldots,s_{k}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $a_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(a_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},a_{k+1},s_{k+1})>0$. Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$. Let $\Omega({\mathcal{M}},\mu)$ denote the subset of $\Omega({\mathcal{M}})$ that includes exactly all runs $r$ such that for every $p\in\Pi({\mathcal{M}})$ that is a prefix of $r$, it holds that $p\in\Pi({\mathcal{M}},\mu)$. Let ${\mathcal{F}}({\mathcal{M}})$ be the set of all strategies.
@@
Given a strategy $\mu$, we can obtain from $\mathcal{M}$} and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\left\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\right\rangle$ in which the choice of actions is resolved according to $\mu$. Formally, for all $p\in\Pi({\mathcal{M}},\mu)$ and $s_{1},s_{2}\in S$ such that both $p_{1}=p\cdot s_{1}$ and $p_{2}=p_{1}\cdot s_{2}$ belong to $\Pi({\mathcal{M}},\mu)$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p_{1},p_{2})=\sum_{a\in A_{s_{1}}}\mu(p_{1})(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s_{1},a,s_{2})$. It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p_{1}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s_{1}$ and $s_{2}$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC. Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state and let $d_{s}\in\mathrm{dist}({\mathcal{A}})$ be the unique distribution such that $\mu(p)=d_{s}$ for all $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$. Note that if no such $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$ exists, then we arbitrarily take $d_{s}$ to be some distribution in $A_{s}$. Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S\times S\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{a\in A_{s}}d_{s}(a)\cdot\mathrm{Pr}_{\mathcal{M}}(s,a,s^{\prime})$. Then, if $p_{0},p_{1},p_{2},\ldots\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then it is not hard to see that the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S$. In other words, the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots$ is sampled according to the finite markov chain $\left\langle S,s_{0},{\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}\right\rangle$.
@@
Let ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\right\rangle$ be an MDP and ${\mathcal{N}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a nondeterministic automaton. When the player takes actions in $\mathcal{M}$} the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of $\mathcal{N}$}. This process can be modeled by an MDP that is a product of $\mathcal{M}$} with $\mathcal{N}$}, and in each step the player takes an action of $\mathcal{M}$} and chooses a successor state in $\mathcal{N}$} that extends the current run on the new letter drawn by $\mathcal{M}$}. Formally, we define the MDP ${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows. The set of states of ${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$. The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows. For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle a,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$. Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle a,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle a,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the $\mathcal{N}$}-run and $\mathcal{M}$}-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{*}\cup(Q_{\mathcal{N}})^{\omega}$ of appropriate lengths, let $r_{\mathcal{M}}\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{N}}\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, we have $\pi_{\mathcal{N}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{N}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{M}}$.

### v6

@@ 6.1. Good-for-MDP automata
We first define MDPs and good-for-MDP automata. A $\Sigma$-labeled Markov decision process (MDP) is ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, where $S$ is a finite set of states, $s_{0}\in S$ is an initial state, and for every $s\in S$, the set $A_{s}$ is a finite set of actions that are available in $s$. Let $A=\bigcup_{s\in S}A_{s}$ be the set of all actions. Then, $\mathrm{Pr}_{\mathcal{M}}:S\times A\times S\nrightarrow[0,1]$ is a (partial) stochastic transition function: for every two states $s,s^{\prime}\in S$ and action $c\in A_{s}$, we have that $\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})$ is the probability of moving from $s$ to $s^{\prime}$ when action $c$ is taken. Accordingly, for every $s\in S$ and $c\in A_{s}$, we have $\Sigma_{s^{\prime}\in S}\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})=1$. Finally, $\Sigma$ is a finite alphabet, and $\tau:S\rightarrow\Sigma$ is a labeling function on the states.
@@
Consider a nondeterministic automaton $\mathcal{N}$} with alphabet $\Sigma$, and a $\Sigma$-labeled MDP $\mathcal{M}$}, and consider the following game between a player and nature. The positions of the game are pairs $\left\langle s,q\right\rangle$, where $s\in S$ is a state of $\mathcal{M}$}, and $q\in Q$ is a state of $\mathcal{N}$}. The game starts by the player picking an action $c\in A_{s_{0}}$ available from the initial state of $\mathcal{M}$}, and the game moves to $\left\langle s,q_{0}\right\rangle$ with probability $\mathrm{Pr}_{\mathcal{M}}(s_{0},c,s)$. Then, at each round from position $\left\langle s,q\right\rangle$, the player chooses an action $c\in A_{s}$ and a state $q^{\prime}\in\delta(q,\tau(s))$, and then nature chooses a state $s^{\prime}\in S$ with probability $\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})$, and the game moves to $\left\langle s^{\prime},q^{\prime}\right\rangle$. The objective of the player is to on the one hand maximize the probability of generating a word in $L({\mathcal{N}})$ and on the other hand, to generate a rejecting run on words in $L({\mathcal{N}})$ with zero probability. When the automaton $\mathcal{N}$} is deterministic, the player has to only pick actions, as the state $q^{\prime}\in\delta(q,\tau(s))$ is unique. When the automaton is nondeterministic, the player may fail to generate an accepting run of $\mathcal{N}$} even when the actions he provides induce a word in $L({\mathcal{N}})$. Indeed, each nondeterministic choice may lead to accepting only a subset of the words in $L({\mathcal{N}})$. Intuitively, $\mathcal{N}$} is good-for-MDP, if for every MDP $\mathcal{M}$}, the nondeterminism in $\mathcal{N}$} does not prevent the player from maximizing the probability of generating a word in $L({\mathcal{N}})$ while generating a rejecting run on words in $L({\mathcal{N}})$ with zero probability. We now formalize this intuition.
@@
A run of $\mathcal{M}$} is an infinite sequence of states $r=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{\omega}$, such that for all $i\geq 0$ there exists $c\in A_{s_{i}}$ with $\mathrm{Pr}_{\mathcal{M}}(s_{i},c,s_{i+1})>0$. A finite run is a prefix of such a sequence. Let $\Omega({\mathcal{M}})\subseteq S^{\omega}$ denote the set of runs of $\mathcal{M}$} from the initial state $s_{0}$, and let $\Pi({\mathcal{M}})\subseteq S^{*}$ denote the set of finite such runs. For a run $r\in\Omega({\mathcal{M}})$, we define the corresponding labeled run as $\tau(r)=\tau(s_{1}),\tau(s_{2}),\tau(s_{3}),\ldots\in\Sigma^{\omega}$. Note that the label of the first state $s_{0}$ of a run is not part of $\tau(r)$.
@@
Consider the set $A$ of actions. Let $\mathrm{dist}(A)$ be the set of distributions over $A$. A distribution $d\in\mathrm{dist}(A)$ is a point distribution if there exists $c\in A$ such that $d(c)=1$. A strategy for an MDP $\mathcal{M}$} is a function $\mu:\Pi({\mathcal{M}})\rightarrow\mathrm{dist}(A)$, which suggests to the player a distribution on the available actions given the history of the game so far. The strategy should suggest an available action, thus $\mu(s_{0},s_{1},s_{2},\ldots,s_{k})(c)>0$ implies that $c\in A_{s_{k}}$. Let $\Pi({\mathcal{M}},\mu)$ denote the subset of $\Pi({\mathcal{M}})$ that includes exactly all finite runs that agree with the strategy $\mu$. Formally, we have that $\left\langle s_{0}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and for all $p^{\prime}=\left\langle s_{0},s_{1},\ldots,s_{k}\right\rangle\in\Pi({\mathcal{M}},\mu)$, and $s_{k+1}\in S$, we have that $p=p^{\prime}\cdot s_{k+1}\in\Pi({\mathcal{M}},\mu)$ iff there exists $c_{k+1}\in A_{s_{k}}$ such that $\mu(p^{\prime})(c_{k+1})>0$ and $\mathrm{Pr}_{{\mathcal{M}}}(s_{k},c_{k+1},s_{k+1})>0$. Thus, a run in $\Pi({\mathcal{M}},\mu)$ can only be extended using actions that have a positive probability according to $\mu$. Let $\Omega({\mathcal{M}},\mu)$ denote the subset of $\Omega({\mathcal{M}})$ that includes exactly all runs $r$ such that for every $p\in\Pi({\mathcal{M}})$ that is a prefix of $r$, it holds that $p\in\Pi({\mathcal{M}},\mu)$. Let ${\mathcal{F}}({\mathcal{M}})$ be the set of all strategies.
@@
Given a strategy $\mu$, we can obtain from $\mathcal{M}$} and $\mu$ the MC ${\mathcal{C}}({\mathcal{M}},\mu)=\left\langle\Pi({\mathcal{M}},\mu),s_{0},\mathrm{Pr}_{\mathcal{M}}^{\mu}\right\rangle$ in which the choice of actions is resolved according to $\mu$. Formally, for all $p\in\Pi({\mathcal{M}},\mu)$ and $s_{1},s_{2}\in S$ such that both $p_{1}=p\cdot s_{1}$ and $p_{2}=p_{1}\cdot s_{2}$ belong to $\Pi({\mathcal{M}},\mu)$, then $\mathrm{Pr}_{\mathcal{M}}^{\mu}(p_{1},p_{2})=\sum_{c\in A_{s_{1}}}\mu(p_{1})(c)\cdot\mathrm{Pr}_{\mathcal{M}}(s_{1},c,s_{2})$. It is clear that when $\mu$ is memoryless, that is, $\mu$ only depends on the last position in $p_{1}$, then ${\mathrm{Pr}}_{\mathcal{M}}^{\mu}$ only depends on $s_{1}$ and $s_{2}$, and hence the projection of the chain ${\mathcal{C}}({\mathcal{M}},\mu)$ onto the last position in a path is a finite MC. Formally, let $\iota:\Pi({\mathcal{M}},\mu)\rightarrow S$ be the function that maps a finite path to its last state and let $d_{s}\in\mathrm{dist}({\mathcal{A}})$ be the unique distribution such that $\mu(p)=d_{s}$ for all $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$. Note that if no such $p\in\Pi({\mathcal{M}},\mu)$ with $\iota(p)=s$ exists, then we arbitrarily take $d_{s}$ to be some distribution in $A_{s}$. Let ${\mathrm{Pr}^{\prime}}_{{\mathcal{M}}}^{\mu}:S\times S\rightarrow[0,1]$ be the stochastic transition function that is defined by ${\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})=\sum_{c\in A_{s}}d_{s}(c)\cdot\mathrm{Pr}_{\mathcal{M}}(s,c,s^{\prime})$. Then, if $p_{0},p_{1},p_{2},\ldots\in(\Pi({\mathcal{M}},\mu))^{\omega}$ is an infinite sequence of increasing paths that are sampled according to ${\mathcal{C}}({\mathcal{M}},\mu)$, then it is not hard to see that the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots\in S^{\omega}$ satisfies $\mathbb{P}(\iota(p_{i+1})=s^{\prime}|\iota(p_{i})=s)={\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}(s,s^{\prime})$ for all $s^{\prime},s\in S$. In other words, the sequence $\iota(p_{0}),\iota(p_{1}),\iota(p_{2}),\ldots$ is sampled according to the finite markov chain $\left\langle S,s_{0},{\mathrm{Pr}^{\prime}}_{\mathcal{M}}^{\mu}\right\rangle$.
@@
Let ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\right\rangle$ be an MDP and ${\mathcal{N}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a nondeterministic automaton. When the player takes actions in $\mathcal{M}$} the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of $\mathcal{N}$}. This process can be modeled by an MDP that is a product of $\mathcal{M}$} with $\mathcal{N}$}, and in each step the player takes an action of $\mathcal{M}$} and chooses a successor state in $\mathcal{N}$} that extends the current run on the new letter drawn by $\mathcal{M}$}. Formally, we define the MDP ${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows. The set of states of ${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$. The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows. For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle c,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$. Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle c,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle c,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,c,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the $\mathcal{N}$}-run and $\mathcal{M}$}-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{*}\cup(Q_{\mathcal{N}})^{\omega}$ of appropriate lengths, let $r_{\mathcal{M}}\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{N}}\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, we have $\pi_{\mathcal{N}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{N}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{M}}$.